Owl Shift 11/25/99

Sasha Glazov (glazov@hep.uchicago.edu)
Thu, 25 Nov 1999 07:04:35 -0600


Shift Summary for owl shift, 25 Nov 99.

On Shift: MT, SG

Hours of beam available at all intensities: 5.30 min
Hours of beam available at nominal intensity: 5.30 min
Hours of E799 data (special runs not included): 5.25 min
Hours of special runs (laser scans, muon runs...): 0 min

Accounting of lost time: Kfrends, pipeline init

========================================================================
Runs ended this shift:
Run 15052 -- E799 run

Tapes brought to FCC this shift:
(including tapes for online split)

none
========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:

Smooth 799 datataking. Accelerator problems for 1h30min

> ========================================================================
Description of Shift:



run spill time comment
------------------------------------------------------------------------
15052 12:00 Start of the shift, run has been running

111-119 2:30 Many beeps from SA4/NM4SEM beacase of slight
overlimit. Checked SA4 ADC, trigger TDC dist. -->
all is fine.

2:50 Disovered that DC4 alignment as measured with
the muon run 15050 is off by ~200 uk. Checked
that it was fine for previous mu-run (15039)

4:55 Beam goes away: Linac klystron station trips off.
Stop the run. Install new DC alignment & pedestals
Estimate for beam -- 1 hour. Start new run

15053 49 6:25 MCR calls and beam comes back in the same minute

========================================================================
Follow up Notes on Previous Shift

---
 
 ========================================================================